# MDDM: A Molecular Dynamics Diffusion Model to Predict Particle Self-Assembly

### v1

@@ 2.1 Dataset
We performed simulations across the parameter ranges: $15.01.0\leq k<$ 15.0, $6.00.0\leq\phi<$ 6.0, $0.050.01\leq T<$ 0.05, with 10 values across each input parameter, for a total of 1000 initial simulations. Visualizations of a few of these results are shown in Figure A1 in the Appendix.
@@ 4.1 Benchmark
Due to of the lack of generative models for self-assembly in the MD domain, we compare against an unorthodox baseline: Diffusion with Transolver. That is, we use Transolver [18] as a means of iteratively denoising a point cloud from noise to structure. The node-embedding capabilities of Transolver make it ideal for processing an input shape and predicting a suitable noise vector at each node. We apply the same diffusion strategy for the baseline as we do for our proposed model, which was inspired in part by Diffusion Point Cloud [13].
@@ 4.3 Training and Sampling
in which $\mathbf{x}_{0}$ is the data, $\mathbf{x}_{t}$ is the data corrupted by $t$ steps of diffusion, and $\alpha_{t}$ is a constant indicating the magnitude of diffusion from step 00 to step $t$. However, use of this equation is predicated on the notion that at the final diffusion step, pixel values are normally distributed. For our systems, due to the presence of PBCs, particle locations will not obey a normal distribution after diffusion. Rather, they will take on uniform distribution, since wrapping a normal distribution makes it become uniform. A proof of this statement can be found in the Appendix. Because the final noise is not normally distributed, the standard DDPM variational inference algorithm must be slightly reformulated. To do so, we redefine the forward diffusion process as follows:
@@
Now, we can continue to use properties of Gaussians, and maximization the Evidence Lower Bound (ELBO) during training can be re-derived for this revised formulation. Note that the result in Eqn. 3 is also wrapped within the periodic box domain. This gives rise to the training and sampling pseudocode below in Algorithms 1 and 2. A probabilistic derivation is provided in the Appendix.
@@ 6 Conclusion
Appendix Appendix C Theorems and Proofs
@@ Appendix C.1 Uniformity of a wrapped Gaussian
Theorem 1.
A normal distribution with standard deviation $L$ that has been wrapped by a periodic boundary function onto the interval $[0,L)$ is equivalent to a uniform distribution on $[0,L)$.
@@
Let $X$ be a random variable with a general Gaussian distribution:
$X$ $\sim N(\mu,L),\quad f_{X}(x)=\frac{1}{\sqrt{2\pi L^{2}}}e^{-\frac{(x-\mu)^{2}}{2L^{2}}}.$
@@
Note that this sum accounts for all possible preimages $\forall k\in\mathds{Z}:\;x=y+kL$, but $F_{Y}(y)$ is only nonzero on the interval $[0,L)$. Also, $|g_{L}^{\prime}(x)|$ evaluates to 1 everywhere. Thus, the sum is:
$f_{Y}(y)=\begin{cases}\frac{1}{L\sqrt{2\pi}}\sum_{k\in\mathds{Z}}e^{-\frac{(y+kL-\mu)^{2}}{2L^{2}}}&\text{if }0\leq y<L,\\ 0&\text{otherwise}\end{cases}$
The sum $S(y,\mu,L)=\sum_{k\in\mathds{Z}}e^{-\frac{(y+kL-\mu)^{2}}{2L^{2}}}$ can be recognized as an instance of the Jacobi theta function $S(y,\mu,L)=\vartheta_{3}(0,1/\sqrt{e})$ [20], which evaluates to a value of $\sqrt{2\pi}$ regardless of $y$, $\mu$, or $L$. Therefore, the PDF simplifies to:
$f_{Y}(y)=\begin{cases}\frac{1}{L}&\text{if }0\leq y<L,\\ 0&\text{otherwise}\end{cases}$
This is trivially recognized as a uniform distribution on $[0,L)$. Therefore, the distribution $\mathcal{N}(\mu,L)$ wrapped onto the interval $[0,L)$ is the uniform distribution $\mathcal{U}[0,L)$.

### v2

@@ 2.1 Dataset
We performed simulations across the parameter ranges: $1.0\leq k\leq 15.0$;  $0.0\leq\phi\leq 6.0$;  $0.01\leq T\leq 0.05$, with 10 values across each input parameter, for a total of 1000 initial simulations. Visualizations of a few of these results are shown in Figure A1 in the Appendix.
@@ 4.1 Benchmark
Due to of the lack of generative models for self-assembly in the MD domain, we compare against a somewhat disparate baseline: Diffusion with Transolver. That is, we use Transolver [18] as a means of iteratively denoising a point cloud from noise to structure. The node-embedding capabilities of Transolver make it ideal for processing an input shape and predicting a suitable noise vector at each node. We apply the same diffusion strategy for the baseline as we do for our proposed model, which was inspired in part by Diffusion Point Cloud [13].
@@ 4.2 Model Architecture
Our network consists of 8 PBC-Conv layers with hidden local/global feature counts of 32 at each layer, using a periodic $k$-nearest neighbors graph with $k=32$. Each PBC-Conv layer contains an MLP with 2 hidden layers of 32 neurons each. The output MLP has 2 layers, each with 128 neurons. The sole global input was $t/T$, current diffusion time step divided by the total number of diffusion steps, with $T=500$ for our experiments.
@@ 4.3 Training and Sampling
in which $\mathbf{x}_{0}$ is the data, $\mathbf{x}_{t}$ is the data corrupted by $t$ steps of diffusion, and $\alpha_{t}$ is a constant indicating the magnitude of diffusion from step 00 to step $t$. However, use of this equation is predicated on the notion that at the final diffusion step, pixel values are normally distributed. For our systems, due to the presence of PBCs, particle locations will not obey a normal distribution after diffusion. Rather, they will take on a (very nearly) uniform distribution, since wrapping a normal distribution makes it become approximately uniform. A proof of this statement can be found in the Appendix. Because the final noise is not normally distributed, the standard DDPM variational inference algorithm must be slightly reformulated. To do so, we redefine the forward diffusion process as follows:
@@
Now, we can continue to use properties of Gaussians, and maximization of the Evidence Lower Bound (ELBO) during training can be re-derived for this revised formulation. Note that the result in Eqn. 3 is also wrapped within the periodic box domain. This gives rise to the training and sampling pseudocode below in Algorithms 1 and 2. A probabilistic derivation is provided in the Appendix.
@@
For our experiments, we use $T=500$ diffusion steps with $\alpha$ on a cosine-schedule $\alpha_{t}=\cos^{2}\left(\frac{\pi}{2}\cdot\frac{x/(T+1)+s}{1+s}\right)$, with $s=0.008$. We train for 800 epochs using an Adam optimizer with learning rate 0.005, which decreases by a factor of 0.95 every 100 epochs.
@@ 6 Conclusion
Acknowledgments
We thank the following individuals for their thoughtful discussion: Thomas O’Connor, Michael Bockstaller, Ayesha Abdullah, Andrew Gillman, Eric Harper, Jeffrey Ethier, Larry Drummy, Jacob Rast, and Matt Gormley.
@@ Acknowledgments
Appendix Appendix C Mathematical claims
@@ Appendix C.1 Uniformity of a wrapped Gaussian
Claim 1.
A normal distribution with standard deviation $L$ that has been wrapped by a periodic boundary function onto the interval $[0,L)$ approximates a uniform distribution on $[0,L)$.
@@
(Note: This proof does not quantify the error in the approximate result. However, our experiments have shown that the probability density function of a wrapped standard normal distribution is virtually indistinguishable from that of a uniform distribution at all points; the following proof explains why.)
We start by showing that the theorem holds for the standard normal distribution with mean 0 and standard deviation 1, and without loss of generality we assert that the result extends to other means and standard deviations.
Let $X$ be a random variable with a standard normal distribution. According to the Central Limit Theorem, this can be closely approximated by an Irwin-Hall distribution with $N=12$ (that is, the distribution describing the sum of 12 independent uniform random variables on $[0,1)$, subtracting 6 here to re-center at zero), which has mean 0 and standard deviation 1 when expressed using the PDF below, adapted from Marengo et al. [20].
$X$ $\overset{\cdot}{\sim}N(0,1),\quad\frac{1}{\sqrt{2\pi}}e^{-\frac{x^{2}}{2}}\approx f_{X}(x)=\frac{1}{2(12-1)!}\sum_{r=0}^{12}(-1)^{r}\binom{12}{r}\text{sgn}(x+6-r)(x+6-r)^{12-1}$
@@
Note that this sum accounts for all possible preimages $\forall k\in\mathds{Z}:\;x=y+kL$, but $f_{Y}(y)$ is only nonzero on the interval $[0,1)$, and only $k$ values from -6 to 5 result in nonzero probability density. Also, $|g_{L}^{\prime}(x)|$ evaluates to 1 everywhere. Thus, the sum is:
$f_{Y}(y)=\begin{cases}{\sum_{k=-6}^{5}\frac{1}{2(12-1)!}\sum_{r=0}^{12}(-1)^{r}\binom{12}{r}\text{sgn}(y+k+6-r)(y+k+6-r)^{12-1}}&\text{if }0\leq y<1,\\ 0&\text{otherwise}\end{cases}$
For y values on the interval $[0,1)$, this expands to:
$f_{Y}(y)\Big{|}_{0\leq y<1}$ $=\Big{(}-462\left(y-6\right)^{11}+\left(y-1\right)^{11}+11\left(y+1\right)^{11}-55\left(y+2\right)^{11}+165\left(y+3\right)^{11}$
$\qquad+330\left(y-5\right)^{11}-55\left(y-10\right)^{11}-11\left(y-2\right)^{11}+11\left(y-11\right)^{11}+55\left(y-3\right)^{11}$
$\qquad-165\left(y+8\right)^{11}+55\left(y+9\right)^{11}-11\left(y+10\right)^{11}+\left(y+11\right)^{11}+462\left(y-7\right)^{11}$
$\qquad-330\left(y-8\right)^{11}-165\left(y-4\right)^{11}+165\left(y-9\right)^{11}-330\left(y+4\right)^{11}+462\left(y+5\right)^{11}$
$\qquad-462\left(y+6\right)^{11}+330\left(y+7\right)^{11}-\left(y-12\right)^{11}-y^{11}\Big{)}\Big{/}\big{(}2\;(12-1)!\big{)}$
$=1$
Through symbolic manipulation, the above expression evaluates to exactly 1. Therefore, the PDF is simply:
$f_{Y}(y)=\begin{cases}1&\text{if }0\leq y<1,\\ 0&\text{otherwise}\end{cases}$
This is recognized as a uniform distribution on $[0,1)$. Therefore, because the wrapped Irwin-Hall distribution is exactly uniform, we conclude the distribution $\mathcal{N}(0,1)$ wrapped onto the interval $[0,1)$ closely approximates the uniform distribution $\mathcal{U}[0,1)$.
Without loss of generality, this relationship holds for any mean $\mu$ and any standard deviation $L$ by rescaling the distributions and shifting the periodic region. Thus, the distribution $\mathcal{N}(\mu,L)$ wrapped onto the interval $[0,L)$ approximates the uniform distribution $\mathcal{U}[0,L)$.